OpenStudy (anonymous):

Convert the mixed number to an improper fraction. 6 7/8

OpenStudy (confluxepic):

\[6\frac{ 7 }{ 8 }=\frac{ 6*8 }{ 8 }+\frac{ 7 }{ 8 }=\frac{ 48+7 }{ 8 }=\frac{ 55 }{ 8 }\]

OpenStudy (confluxepic):

\[\frac{ 55 }{ 8 }\]